# rem $\diamond$

### v1

@@ 4 Connections to Model Theory
$h\colon\mathbb{R}^{n}\to\{0,1\},\ \underline{x}\mapsto\mathds{1}_{[0,\infty)}\!\left(\sum_{i=1}^{k}u_{i}\sigma(\underline{v}_{i}\cdot\underline{x}+\underline{v}_{i,0})+u_{0}\right)\!,$
where $\underline{x}\in\mathbb{R}^{n}$ is the input vector, $u_{i}\in\mathbb{R}$ (for $i\in[k]_{0}$) are the output weights, $\underline{v}_{i},\underline{v}_{i,0}\in\mathbb{R}^{n}$ (for $i\in[k]$) are the input weights, and the activation function $\sigma\colon\mathbb{R}\to\mathbb{R}$ is the standard sigmoid function
$\sigma(x)=\frac{1}{1+\exp(x)}.$
@@
Given a language $\mathcal{L}$ expanding $\mathcal{L}_{<}$, and an $\mathcal{L}$-structure $\mathcal{M}$ whose reduct $(M,<)$ is a linear order, recall that $\mathcal{M}$ is o-minimal if any set $A\subseteq M$ that is definable over $\mathcal{M}$ can be expressed as a finite union of points and open intervals (see Marker [18, Definition 3.1.18]). Steinhorn [29, page 27] notes that “PAC learning is directly relevant to neural networks whose architecture defines concept classes that consist of definable families of definable sets in o-minimal structures.” Most crucially, due to Wilkie [33] the real exponential field $\mathbb{R}_{\exp}$ is an o-minimal $\mathcal{L}_{\exp}$–structure. O-minimality implies the property NIP (see Pillay and Steinhorn [21, Corollary 3.10]), which in turn ensures that hypothesis spaces of the form $\mathcal{H}_{\mathcal{X}}^{\varphi}$ have finite VC dimension. All hypothesis spaces as in Example 4.2 are therefore of finite VC dimension. As mentioned in the introduction, these observations were first made by Laskowski [16]. Since Laskowski’s original arguments dealt with quantitative bounds for concept classes rather than hypothesis spaces over definable instance spaces $\mathcal{X}$, we present here a self-contained and more direct proof of these observations. We first introduce the relevant model-theoretic property NIP (see Poizat [22, § 12.4]), which is originally due to Shelah [26].
@@
Definition 4.3 expresses that a formula $\varphi(\underline{x};\underline{p})$ has NIP if there is $m\in\mathbb{N}$ such that no set of size $m$ is shattered by $\varphi$ (see Simon [28, Remark 2.3]).
@@
Proposition 4.5 closely relates the model-theoretic notion NIP with the concept of VC dimension from Statistical Learning Theory. Namely, any hypothesis space $\mathcal{H}^{\varphi}_{\mathcal{X}}$ over an NIP structure has finite VC dimension. On the other hand, the Fundamental Theorem of Statistical Learning connects VC dimension with PAC learning. Therefore, drawing a direct connection between NIP and PAC learning suggests itself (cf. e.g. Steinhorn [29, page 27]). However, the measurability requirements that we carefully scrutinized in Section 3 have to be accounted for.
@@
Let $\mathcal{L}$ be a language expanding $\mathcal{L}_{\mathrm{or}}$, let $\mathcal{R}$ be an o-minimal $\mathcal{L}$–expansion of $\mathbb{R}_{\mathrm{or}}$, let $n,\ell\in\mathbb{N}$, let $\varphi(x_{1},\ldots,x_{n};p_{1},\dots,p_{\ell})$ be an $\mathcal{L}$–formula and let $\underline{w}\in\mathbb{R}^{\ell}$. Then $\varphi(\mathcal{R};\underline{w})\in\mathcal{B}(\mathbb{R}^{n})$.
@@
Due to the Cell Decomposition Theorem (van den Dries [9, page 52]), the set $\varphi(\mathcal{R};\underline{w})$ has a finite partition into cells (see Definition A.8). It is readily verified that cells are Borel sets (see Lemma A.9), and hence $\varphi(\mathcal{R};\underline{w})\in\mathcal{B}(\mathbb{R}^{n})$. ∎
@@
Thus, the map $\hat{\operatorname{er}}_{\square}(h_{\square})$ is $(\Sigma_{\mathcal{Z}}^{m}\otimes\mathcal{B}(\mathbb{R}^{\ell}))$–measurable. The measurability of the maps $\operatorname{er}_{\mathbb{D}}(h_{\square})$ and $\hat{\operatorname{er}}_{\square}(_{\square})$ implies that the maps $f_{U}$ and $f_{V}$ are measurable. Since we assume that the probability space $(\mathcal{Z}^{k},\Sigma_{\mathcal{Z}}^{k},\mathbb{D}^{k})$ is complete for any $k\in\mathbb{N}$, we can apply the arguments of Pollard [23, page 197] to deduce that the map
@@ A.1 Auxiliary Lemmas
The following result is a refined version of Shalev-Shwartz and Ben-David [25, Lemma A.4].
@@
If $\beta\geq e$, then applying [25, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=\beta$) yields
@@
Otherwise, applying [25, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=e$) yields
@@ A.2 Well-Behaved Hypothesis Spaces
is $\Sigma_{\mathcal{Z}}^{2m}$–measurable. In general, we however need to take suprema over uncountable families of these measurable functions. In this situation, verifying the measurability of the maps $U$ and $V$ is more complicated or even impossible (see Example A.13). However, in this section we aim at identifying further sufficient conditions that might be helpful in the uncountable setting as well.
We now introduce the concept of a universally separable hypothesis space. Dudley [10, page 902] was the first to introduce this property “as a way of avoiding measurability difficulties” (Pollard [23, page 38]).
@@
We next prove that universal separability ensures measurability of the maps $U$ and $V$ from Definition 3.2. Similar statements in a more general setting can be found in the literature (cf. e.g. Ben-David, Mansour and Benedek [3, Lemma 5.2] and Pollard [23, Chapter II, Problem 3]). Guided by the proof of [3, Lemma 5.2] and by arguments suggested by D. Pollard [^11] The second author thanks D. Pollard for these suggestions via private communication., we present here a detailed proof adapted to the maps $U$ and $V$.
@@
The universal separability condition applies to many standard examples of hypothesis spaces (cf. Pollard [23, Chapter II, Problems 4, 5 and 7]). However, universal separability is no necessary condition for ensuring well-behavedness. In fact, Example A.15 presents a quite simple hypothesis space that is not universally separable, but well-behaved with respect to certain sets of distributions.
Further considerations on the measurability of suprema over uncountable families of measurable functions can be found in Dudley [10], [11, Chapter 10], Gaenssler [12, Chapter 4] and Pollard [23, Appendix C]. For instance, [23, Appendix C] introduces the concept of permissible function classes, which Haussler [14, Section 9.2] also takes up. Moreover, recall that the arguments in [23, Appendix C] constitute the core for our proof of Theorem 4.7.
@@ A.3 Cells
In this section we mostly follow van den Driess [9, Chapter 3, § 2]. Let $\mathcal{L}$ be a language expanding $\mathcal{L}_{<}$ and let $\mathcal{R}$ be an o-minimal $\mathcal{L}$–structure. We endow the domain $R$ with the order-topology $\tau$. Given $m\in\mathbb{N}$ and a set $X\subseteq R^{m}$, we endow $R^{m}$ with the product topology $\tau^{m}$ and $X$ with the subset topology $\tau_{X}=\{\mathcal{O}\cap X\mid\mathcal{O}\in\tau^{m}\}$. Recall that $(R^{m},\tau^{m})$ is a Hausdorff topological space. We denote by $\mathcal{B}(R^{m})$ the Borel $\sigma$–algebra on $R^{m}$, which is the smallest $\sigma$–algebra containing all open sets in $R^{m}$. Further, we denote by $\mathcal{B}(X)$ the trace $\sigma$–algebra given by $\mathcal{B}(X)=\{B\cap X\mid B\in\mathcal{B}(R^{m})\}$. Recall that $\mathcal{B}(X)$ is the smallest $\sigma$–algebra containing $\tau_{X}$, and $\mathcal{B}(X)=\{B\in\mathcal{B}(R^{m})\mid B\subseteq X\}\subseteq\mathcal{B}(R^{m})$ if and only if $X\in\mathcal{B}(R^{m})$ (see Bogachev [7, Lemma 6.2.4]). A map $f\colon X\to R$ is called definable if its graph $\Gamma(f)=\{(\underline{x},f(\underline{x}))\mid\underline{x}\in X\}\subseteq R^{m+1}$ is definable, and it is called continuous if the preimage of any open set $\mathcal{O}\in\tau$ is contained in $\tau_{X}$. If $X$ is definable, then we set

### v2

@@ 4 Connections to Model Theory
$h\colon\mathbb{R}^{n}\to\{0,1\},\ \underline{x}\mapsto\mathds{1}_{[0,\infty)}\!\left(\sum_{i=1}^{k}u_{i}\sigma(\underline{v}_{i}\cdot\underline{x}+v_{i,0})+u_{0}\right)\!,$
where $\underline{x}\in\mathbb{R}^{n}$ is the input vector, $u_{i}\in\mathbb{R}$ (for $i\in[k]_{0}$) are the output weights, $\underline{v}_{i}\in\mathbb{R}^{n},v_{i,0}\in\mathbb{R}$ (for $i\in[k]$) are the input weights, and the activation function $\sigma\colon\mathbb{R}\to\mathbb{R}$ is the standard sigmoid function
$\sigma(x)=\frac{1}{1+\exp(-x)}.$
@@
More generally, if an artificial neural network for binary classification only uses activation functions that are parameter-free definable over $\mathbb{R}_{\exp}$ (i.e. their graphs are definable over $\mathbb{R}_{\exp}$ by a formula without parameters), then the same as above holds. Namely, for any such hypothesis class $\mathcal{H}$, there exists an $\mathcal{L}_{\exp}$–formula $\varphi$ such that over $\mathbb{R}_{\exp}$ we have $\mathcal{H}=\mathcal{H}^{\varphi}$. Many activation functions that are commonly used in practice are parameter-free definable over $\mathbb{R}_{\exp}$. For instance, the graph of the ReLU function $\mathrm{ReLU}(x)=\max\{0,x\}$ is defined by the $\mathcal{L}_{\exp}$–formula
$(x<0\to y=0)\wedge(0\leq x\to y=x).\vspace{-0.7em}$
Given a language $\mathcal{L}$ expanding $\mathcal{L}_{<}$, and an $\mathcal{L}$-structure $\mathcal{M}$ whose reduct $(M,<)$ is a linear order, recall that $\mathcal{M}$ is o-minimal if any set $A\subseteq M$ that is definable over $\mathcal{M}$ can be expressed as a finite union of points and open intervals (see Marker [20, Definition 3.1.18]). Steinhorn [31, page 27] notes that “PAC learning is directly relevant to neural networks whose architecture defines concept classes that consist of definable families of definable sets in o-minimal structures.” Most crucially, due to Wilkie [35] the real exponential field $\mathbb{R}_{\exp}$ is an o-minimal $\mathcal{L}_{\exp}$–structure. O-minimality implies the property NIP (see Pillay and Steinhorn [23, Corollary 3.10]), which in turn ensures that hypothesis spaces of the form $\mathcal{H}_{\mathcal{X}}^{\varphi}$ have finite VC dimension. All hypothesis spaces as in Example 4.2 are therefore of finite VC dimension. As mentioned in the introduction, these observations were first made by Laskowski [18]. Since Laskowski’s original arguments dealt with quantitative bounds for concept classes rather than hypothesis spaces over definable instance spaces $\mathcal{X}$, we present here a self-contained and more direct proof of these observations. We first introduce the relevant model-theoretic property NIP (see Poizat [24, § 12.4]), which is originally due to Shelah [28].
@@
Definition 4.3 expresses that a formula $\varphi(\underline{x};\underline{p})$ has NIP if there is $m\in\mathbb{N}$ such that no set of size $m$ is shattered by $\varphi$ (see Simon [30, Remark 2.3]).
@@
Proposition 4.5 closely relates the model-theoretic notion NIP with the concept of VC dimension from Statistical Learning Theory. Namely, any hypothesis space $\mathcal{H}^{\varphi}_{\mathcal{X}}$ over an NIP structure has finite VC dimension. On the other hand, the Fundamental Theorem of Statistical Learning connects VC dimension with PAC learning. Therefore, drawing a direct connection between NIP and PAC learning suggests itself (cf. e.g. Steinhorn [31, page 27]). However, the measurability requirements that we carefully scrutinized in Section 3 have to be accounted for.
@@
Let $\mathcal{L}$ be a language expanding $\mathcal{L}_{\mathrm{or}}$, let $\mathcal{R}$ be an o-minimal $\mathcal{L}$–expansion of $\mathbb{R}_{\mathrm{or}}$, let $n,\ell\in\mathbb{N}$, let $\varphi(x_{1},\ldots,x_{n};p_{1},\dots,p_{\ell})$ be an $\mathcal{L}$–formula and let $\underline{w}\in\mathbb{R}^{\ell}$. Then $\varphi(\mathcal{R};\underline{w})\in\mathcal{B}(\mathbb{R}^{n})$. [^9] In Kaiser [16, Proposition 1.1] a similar result is stated and proved.
@@
Due to the Cell Decomposition Theorem (van den Dries [10, page 52]), the set $\varphi(\mathcal{R};\underline{w})$ has a finite partition into cells (see Definition A.8). It is readily verified that cells are Borel sets (see Lemma A.9), and hence $\varphi(\mathcal{R};\underline{w})\in\mathcal{B}(\mathbb{R}^{n})$. ∎
@@
Thus, the map $\hat{\operatorname{er}}_{\square}(h_{\square})$ is $(\Sigma_{\mathcal{Z}}^{m}\otimes\mathcal{B}(\mathbb{R}^{\ell}))$–measurable. The measurability of the maps $\operatorname{er}_{\mathbb{D}}(h_{\square})$ and $\hat{\operatorname{er}}_{\square}(_{\square})$ implies that the maps $f_{U}$ and $f_{V}$ are measurable. Since we assume that the probability space $(\mathcal{Z}^{k},\Sigma_{\mathcal{Z}}^{k},\mathbb{D}^{k})$ is complete for any $k\in\mathbb{N}$, we can apply the arguments of Pollard [25, page 197] to deduce that the map
@@
  5. (e)
    Theorem 4.7 can potentially be extended to other structures that are, in some sense, tame (e.g. o-minimal structures, structures with quantifier elimination) and naturally endowed with a topology (cf. [17, § 5]). However, one must carefully examine whether the arguments of Pollard [25, page 197] – or those upon which Pollard’s approach relies (cf. Dellacherie and Meyer [9, Chapter III]) – can still be applied in such settings. These arguments, for instance, require the topological space to be metrizable, separable and locally compact.
@@ A.1 Auxiliary Lemmas
The following result is a refined version of Shalev-Shwartz and Ben-David [27, Lemma A.4].
@@
If $\beta\geq e$, then applying [27, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=\beta$) yields
@@
Otherwise, applying [27, Lemma A.4] (with $x^{\prime}=0,a=\alpha,b=e$) yields
@@ A.2 Well-Behaved Hypothesis Spaces
is $\Sigma_{\mathcal{Z}}^{2m}$–measurable.
In general, we however need to take suprema over uncountable families of these measurable functions. In this situation, verifying the measurability of the maps $U$ and $V$ is more complicated or even impossible (see Example A.13). However, in this section we aim at identifying further sufficient conditions that might be helpful in the uncountable setting as well.
We now introduce the concept of a universally separable hypothesis space. Dudley [11, page 902] was the first to introduce this property “as a way of avoiding measurability difficulties” (Pollard [25, page 38]).
@@
We next prove that universal separability ensures measurability of the maps $U$ and $V$ from Definition 3.2. Similar statements in a more general setting can be found in the literature (cf. e.g. Ben-David, Mansour and Benedek [3, Lemma 5.2] and Pollard [25, Chapter II, Problem 3]). Guided by the proof of [3, Lemma 5.2] and by arguments suggested by D. Pollard [^13] The second author thanks D. Pollard for these suggestions via private communication., we present here a detailed proof adapted to the maps $U$ and $V$.
@@
The universal separability condition applies to many standard examples of hypothesis spaces (cf. Pollard [25, Chapter II, Problems 4, 5 and 7]). However, universal separability is no necessary condition for ensuring well-behavedness. In fact, Example A.15 presents a quite simple hypothesis space that is not universally separable, but well-behaved with respect to certain sets of distributions.
Further considerations on the measurability of suprema over uncountable families of measurable functions can be found in Dudley [11], [12, Chapter 10], Gaenssler [13, Chapter 4] and Pollard [25, Appendix C]. For instance, [25, Appendix C] introduces the concept of permissible function classes, which Haussler [15, Section 9.2] also takes up. Moreover, recall that the arguments in [25, Appendix C] constitute the core for our proof of Theorem 4.7.
@@ A.3 Cells
In this section we mostly follow van den Driess [10, Chapter 3, § 2]. Let $\mathcal{L}$ be a language expanding $\mathcal{L}_{<}$ and let $\mathcal{R}$ be an o-minimal $\mathcal{L}$–structure. We endow the domain $R$ with the order-topology $\tau$. Given $m\in\mathbb{N}$ and a set $X\subseteq R^{m}$, we endow $R^{m}$ with the product topology $\tau^{m}$ and $X$ with the subset topology $\tau_{X}=\{\mathcal{O}\cap X\mid\mathcal{O}\in\tau^{m}\}$. Recall that $(R^{m},\tau^{m})$ is a Hausdorff topological space. We denote by $\mathcal{B}(R^{m})$ the Borel $\sigma$–algebra on $R^{m}$, which is the smallest $\sigma$–algebra containing all open sets in $R^{m}$. Further, we denote by $\mathcal{B}(X)$ the trace $\sigma$–algebra given by $\mathcal{B}(X)=\{B\cap X\mid B\in\mathcal{B}(R^{m})\}$. Recall that $\mathcal{B}(X)$ is the smallest $\sigma$–algebra containing $\tau_{X}$, and $\mathcal{B}(X)=\{B\in\mathcal{B}(R^{m})\mid B\subseteq X\}\subseteq\mathcal{B}(R^{m})$ if and only if $X\in\mathcal{B}(R^{m})$ (see Bogachev [7, Lemma 6.2.4]). A map $f\colon X\to R$ is called definable if its graph $\Gamma(f)=\{(\underline{x},f(\underline{x}))\mid\underline{x}\in X\}\subseteq R^{m+1}$ is definable, and it is called continuous if the preimage of any open set $\mathcal{O}\in\tau$ is contained in $\tau_{X}$. If $X$ is definable, then we set
@@ A.4 Examples
Data and Materials Availability: Not applicable.
Code Availability: Not applicable.
Ethical Approval: Not applicable.
Consent to Participate: Not applicable.
Consent for Publication: Not applicable. The paper does not include data or images that require permissions to be published.